Q#4.a. Find the Taylor Series for : f(x) = 7x2 - 6x + 1
Answer #1

Taylor series for fla) about x=a 11 2 3 (x-a) fla) + (x-a) fla) + dex) = f(x) +(2-a) f(a) d! fe-e) = 7 78-62+) fra)= má tá-6a